If the measure of angle a is three-eighths of angle b, and both are supplementary what is the measure of each angle
Answer by tommyt3rd(5050)

a=3/8 b
a+b=180
so...
3/8 b + b = 180
11/8 b = 180
b= 1440/11
a=540/11